If an employee deposits Rs. 2,000 at the end of each year into his company’s plan which pays 7% interest compounded quarterly, how much will he have in the account at the end of 5 years?
1
Expert's answer
2019-05-21T04:26:40-0400

The interest is compounded quarterly, with a year total of 7%. This means that effectively 7/4 = 1.75% a paid every quarter, leading to a year total interest of "1.0175^4 - 1=0.071859"

If the employer starts his first year with "Sum_i = 2000" on his account then on the end of first year he will have


"Sum_1=2000+Sum_i*0.071859+Sum_i=2000+1.071859 *Sum_i=4143.718"


For the year 2 the situation is analogous


"Sum_2=2000+1.071859 *Sum_2 = 2000+1.071859*4143.718= 6441.48"


For the end of year 3 and 4


"Sum_3=2000+1.071859 *Sum_2 = 2000+1.071859*6441.48= 8904.36"


"Sum_4=2000+1.071859 *Sum_3 = 2000+1.071859*8904.36= 11544.22"

and finally an employee will have


"Sum_5=2000+1.071859 *Sum_4 = 2000+1.071859*11544.22= 14373.78"

Rs. in his account.